Hardwig Inc. is considering whether to pursue a restricted or relaxed current asset investment policy. The firm's annual sales are expected to total $3,600,000, its fixed assets turnover ratio equals 4.0, and its debt and common equity are each 50% of total assets. EBIT is $150,000, the interest rate on the firm's debt is 10%, and the tax rate is 40%. If the company follows a restricted policy, its total assets turnover will be 2.5. Under a relaxed policy its total assets turnover will be 2.2.

Refer to the data for Hardwig, Inc. What's the difference in the projected ROEs under the restricted and relaxed policies?

a. 1.20%
b. 1.80%
c. 1.50%
d. 2.16%
e. 2.59%
Answer #1

Fixed assets = Sales/Fixed asset turnover = 3600000/4 = 900,000

1. Calculations in restricted policy -

Asset turnover = Sales/ Total assets = 2.5

Total assets = 3600000/2.5 = 1,440,000

Debt = Equity = 50% of Assets = 0.5 * 1,440,000 = 720,000

EBIT = 150,000

Interest = 10% * 720,000 = 72,000

Profit before tax = 150,000 - 72,000 = 78,000

Tax = 40% * 78,000 = 31,200

Profit after tax = 78,000 - 31,200 = 46,800

ROE = 46,800/ 720,000 = 6.5%

2. Calculations in relaxed policy -

Asset turnover = Sales/ Total assets = 2.2

Total assets = 3600000/2.2 = 1,636,364

Debt = Equity = 50% of Assets = 0.5 * 1,636,364 = 818,182

EBIT = 150,000

Interest = 10% * 818,182 = 81,818

Profit before tax = 150,000 - 81,818 = 68,182

Tax = 40% * 68,182 = 27,273

Profit after tax = 68,182 - 27,273 = 40,909

ROE = 40,909/ 818,182 = 5%

Difference between ROE under the 2 methods = 6.5% - 5% = 1.5%

Final answer : c) 1.5%